# Towards Stealthy and Effective Backdoor Attacks on Lane Detection: A Naturalistic Data Poisoning Approach

### v1

@@ Introduction
We evaluate DBALD to answer the following research questions:
  1. RQ1.
    How effective is DBALD for attacking LD models, compared with the state-of-the-art methods?
  2. RQ2.
    How stealthy are the triggers and poisoned images generated by DBALD, compared to state-of-the-art methods?
  3. RQ3.
    How effective are the technical components of DBALD in an ablation study?
  4. RQ4.
    How effective is DBALD in physical attack scenarios?
@@
[…]
Lane Detection Model. We adopt four widely used LD models from the literature as our target benchmark, including two anchor-based lane detection methods, LaneATT [32] and ADNet [33], and two segmentation-based methods, RESA [34] and SCNN [35]. These models represent a diverse range of architectural designs and detection principles, enabling a comprehensive evaluation landscape.
@@
Attack Baseline. Considering practical deployment in autonomous driving and LD tasks, we adopt four state-of-the-art and physically realizable backdoor attack strategies as our baselines: ① Fixed Pattern - BadNets [56]: Adds a fixed white square at the bottom-right corner of the image as a static trigger pattern. ② Fixed Image - Blended [41]: Blends a predefined universal image trigger with the input sample at a fixed blending ratio. ③ Fixed Pattern & Image - LD-Attack [15]: Uses a fixed trigger overlay and blends it with the input image. ④ Random Texture - BadLane [16]: Introduces randomly placed mud-like texture noise across the road surface, simulating naturally occurring triggers.
@@
Evaluation Metric. We follow standard TuSimple protocol [16], reporting Accuracy (ACC) on clean data and Attack Success Rate (ASR) under LOA, LRA, and LDA. ACC and ASR for LOA/LRA are defined as $\text{Metric}=\frac{\sum_{i}C_{i}}{\sum_{i}S_{i}}$, where $C_{i}$ is the number of correct lane points (within 20px of GT), and $S_{i}$ is the total. For LDA, ASR is computed as the fraction of disappeared lanes.
@@
Implementation. We set the poisoning rate to 10%, with 900-pixel triggers. For BadNets/Blended, a ${\times}30$ 30 square is placed bottom-right. LD-Attack place 900-pixel on the adjacent lane near the boundary. BadLane samples 900 pixels from a ${\times}100$ 100 region. Our method uses a fixed ${\times}40$ 40 mask with diffusion-based trigger generation, using prompts such as “Add a small and sparse brown mud” or “Add a small traffic cone.” UltraEdit [22] is used for generation. For attack parameters, We follow prior work and set the LOA offset to 60 pixels and the LRA rotation to 9 degrees.
RQ1: The Effectiveness of DBALD
@@
We compare different backdoor attack baselines on four LD models across the CULane and TuSimple datasets. Specifically, we evaluate: (1) the accuracy (ACC) of different LD models on benign samples and poisoned samples; (2) the performance of infected LD models on clean test sets, to assess whether backdoor injection affects normal functionality; and (3) the ASR under various backdoor injection strategies including LDA, LRA, and LOA.
@@
The experimental results are given in Table 1, from which we draw the following key observations: i) CULane is more challenging to attack than TuSimple. The CULane dataset consists of complex urban driving scenes with dynamic objects such as pedestrians, traffic lights, and intersections, making it significantly more difficult to manipulate compared to the highway-based TuSimple dataset. Traditional attack methods perform poorly on both datasets, and although Badlane achieves relatively high ASR on TuSimple (e.g., 90.84% on LaneATT), its performance drops significantly on CULane (only 54.56%). We remark that this highlights the difficulty of launching successful physical backdoor attacks in more diverse and cluttered environments; ii) DBALD achieves the highest ASR across all attack strategies. Our proposed method outperforms all baseline methods under multiple attack strategies, including LDA, LRA, and LOA. Among these, LDA consistently yields the highest ASR, likely due to its simplicity in altering lane direction. In contrast, LOA and LRA are harder to execute but pose greater safety risks by shifting the vehicle trajectory. LOA achieves higher ASR than LRA, which may be attributed to a broader label shift. Overall, DBALD improves the average ASR by a substantial margin and generalizes well across different lane detection models. This further demonstrates that the physical trigger is more effective in activating the backdoor, consistent with findings reported in other backdoor attacks targeting autonomous driving systems [57]; iii) DBALD maintains high accuracy on clean samples. Despite its strong attack capability, DBALD does not degrade model performance on clean inputs. Infected models retain comparable accuracy to their vanilla counterparts, suggesting that the injected backdoor remains stealthy and does not interfere with normal behavior. iv) ADNet is more robust than other lane detection models. LaneATT, RESA, and SCNN models exhibit high ASRs under DBALD, nearly matching their clean ACCs, indicating high vulnerability. In contrast, ADNet shows greater resilience, with a notable gap of over 16% between its ASR and ACC, suggesting that model architecture plays a role in robustness against physical backdoor attacks.
@@
RQ2: The Stealthiness of DBALD
To evaluate the stealthiness of the injected backdoor, we utilize existing image-based backdoor detectors, as no dedicated backdoor detectors for LD models are currently available in the literature. Specifically, we employ three representative methods: UniDetection (U) [19], LGrad (L) [20], and DIRE (D) [58]. UniDetection leverages large-scale pretrained models such as CLIP-ViT to extract high-dimensional visual features and capture subtle differences between generated and natural images. LGrad utilizes intermediate feature maps from CNN-based backbones (e.g., ResNet50) to compute gradient-based saliency maps which reveal local visual artifacts indicative of manipulated content. DIRE is specifically designed to detect images synthesized by diffusion models.
We feed both benign samples and poisoned samples into each detector. All three methods achieve over 90% accuracy on benign samples. The detection results for poisoned samples are summarized in Table 3. As shown, BadNets and Blended remain largely undetected, likely due to their trigger being directly overlaid onto the image without introducing detectable artifacts. In contrast, UniDetection detects nearly 60% of Badlane’s backdoored samples, while only identifying less than 3% of DBALD’s samples. Unfortunately, LGrad fails to effectively detect both DBALD and Badlane, possibly because the detectors are not fine-tuned on autonomous driving related scenarios. Additionally, DIRE proves ineffective against DBALD samples. This is expected, as DBALD only modifies a small region of the image while preserving most of the original content. Note that, although BadNets and Blended exhibit a lower detection rate by forensic detectors, they are indeed visually conspicuous and thus vulnerable to manual inspection. As illustrated in Figure 5, both methods introduce clearly distinguishable patterns (e.g., large square patches or blended logos), while the mud trigger from DBALD is visually natural and blends into the road context, making it much harder to identify through manual observation. This highlights the dual stealthiness of DBALD in both automated detection and human perceptual inspection. More trigger visualizations can be found at [53].
Figure 5: Trigger stealthiness in different attack methods
Method CULane TuSimple
U L D U L D
BadNets 2.56 N/A N/A 0.08 N/A N/A
Blended 5.18 N/A N/A 0.17 N/A N/A
LD-Attack 6.27 N/A N/A 1.31 N/A N/A
Badlane 59.31 9.17 N/A 35.60 4.23 N/A
DBALD 2.72 1.76 0.56 0.23 0.51 0.04
RQ3: Ablation Study
Effectiveness of Heatmap Guidance. We investigate the contribution of heatmap guidance in enhancing the attack robustness of DBALD. All experiments are conducted under the mud trigger on the CULane dataset. Results are given in Table 5, where “Without heatmap” means that the mud trigger is placed randomly during poisoning without any heatmap-based guidance. We find that using heatmap guidance significantly improves ASR on both models: +3.21% on LaneATT and +6.77% on RESA. These results demonstrate that spatially guided trigger placement enhances the robustness of the backdoor attack under dynamic attack strategies. Besides, even the randomly placed trigger achieves higher ASR than Badlane, further indicating that physical triggers are more effective than simple image overlay methods.
We further conduct an analytical study by visualizing the gradient-based attention maps over training epochs for both clean and trigger-injected images. We observe that, during training, attention in clean images gradually concentrates on task-relevant semantic regions, whereas in trigger-injected images, it progressively shifts toward the trigger location. This evolving focus suggests that the model gradually “memorizes” the trigger, reinforcing the effectiveness of targeted trigger placement (see Table 4).
Table 4: Comparison of gradient-based attention at epoch 100.
Setting Attention Entropy $\downarrow$
Attention on
Trigger Region $\uparrow$
Trigger (Ours) 16.09 47.39%
Random Trigger 16.26 44.13%
Clean Image 16.96 41.58%
Table 5: Attack performance w/wo the heatmap guidance.
Method LaneATT RESA
LDA LRA LOA LDA LRA LOA
With heatmap 81.65 73.45 74.19 78.15 65.15 76.86
Without heatmap 80.97 70.16 68.52 74.23 58.26 69.37
@@
ASR under different attack parameters We further explore the impact of trigger size under the LOA setting using the LaneATT model and CULane dataset with mud triggers. As shown in Figure 6, smaller triggers yield significantly lower attack success rates (ASR), with ASR dropping from 74.19% (40$\times$40) to 32.15% (10$\times$10). This suggests that tiny triggers are less effective at activating the backdoor. However, ASR slightly drops when the trigger size exceeds ${\times}40$ 40 (e.g., 70.62% at ${\times}50$ 50). This is due to a corresponding drop in clean accuracy—excessively large triggers distort the input scene, degrading overall model performance and thus suppressing backdoor activation. Notably, even with a trigger as small as 20$\times$20 pixels, the ASR exceeds 50%, indicating that DBALD remains effective under compact trigger settings. This supports its potential stealthiness in practical deployment.
Effectiveness of Poisoning Rate. We further study how different poisoning rates affect the attack performance of DBALD. On both LaneATT and RESA models, we train LD models using poisoning rates of 1%, 3%, 5%, 10% and 15% on the CULane dataset. For reference, constructing a 1% poisoned dataset (i.e., 177 poisoned samples) takes approximately 3 minutes. The results are given in Table 6. We find that DBALD achieves relative strong attack performance even at very low poisoning rates. As the poisoning rate rises, the ASR improves gradually, while the accuracy on clean samples decreases slowly.

### v2

@@ Introduction
Datasets. We use CULane [pan2018SCNN] and TuSimple [pizzati2019lane], containing 8888K/3434K and 3.63.6K/2.72.7K train/test images respectively, with resolutions $1,640{\times}590$ and $1,280{\times}720$, and up to 4 or 5 lanes per frame.
Lane Detectors. We evaluate four representative models: anchor-based model (LaneATT [tabelini2021keep], ADNet [xiao2023adnet]) and segmentation-based model (SCNN [pan2018spatial] RESA [zheng2021resa]).
@@
Attack Baseline. Considering practical deployment in autonomous driving and LD tasks, we adopt four state-of-the-art and physically realizable backdoor attack strategies as our baselines: ① Fixed Pattern - BadNets [gu2019badnets]: Adds a fixed white square at the bottom-right corner of the image as a static trigger pattern. ② Fixed Image - Blended [chen2017targeted]: Blends a predefined universal image trigger with the input sample at a fixed blending ratio. ③ Fixed Pattern & Image - LD-Attack [han2022physical]: Uses a fixed trigger overlay and blends it with the input image. ④ Random Texture - BadLane [zhang2024towards]: Introduces randomly placed mud-like texture noise across the road surface, simulating naturally occurring triggers.
@@
Evaluation Metric. We follow standard TuSimple protocol [zhang2024towards], reporting Accuracy (ACC) on clean data and Attack Success Rate (ASR) under LOA, LRA, and LDA. ACC and ASR for LOA/LRA are defined as $\text{Metric}=\frac{\sum_{i}C_{i}}{\sum_{i}S_{i}}$, where $C_{i}$ is the number of correct lane points (within 20px of GT), and $S_{i}$ is the total. For LDA, ASR is computed as the fraction of disappeared lanes.
@@
Implementation. We set the poisoning rate to 10%, with 900-pixel triggers. For BadNets/Blended, a $30{\times}30$ square is placed bottom-right. LD-Attack place 900-pixel on the adjacent lane near the boundary. BadLane samples 900 pixels from a $100{\times}100$ region. Our method uses a $40{\times}40$ mask with diffusion-based trigger generation, using prompts such as “Add a small and sparse brown mud” or “Add a small traffic cone”. UltraEdit [zhao2024ultraedit] is used for generation. For attack parameters, we follow prior work and set the LOA offset to 60 pixels and the LRA rotation to 9 degrees.
The Effectiveness of DBALD
@@
We compare different backdoor attack baselines on four LD models across the CULane and TuSimple datasets. Specifically, we evaluate: (1) the accuracy (ACC) of original/benign and poisoned LD models on the clean images to assess the clean accuracy drop of different attack methods; and (2) the ASR under various backdoor injection strategies, including LDA, LRA, and LOA.
@@
The experimental results tested on the mud trigger are given in Table 1, from which we draw the following key observations: i) CULane is more challenging to attack than TuSimple. The CULane dataset consists of complex urban driving scenes with dynamic objects such as pedestrians, traffic lights, and intersections, making it significantly more difficult to manipulate compared to the highway-based TuSimple dataset. Traditional attack methods perform poorly on both datasets, and although Badlane achieves relatively high ASR on TuSimple (e.g., 90.84% on LaneATT), its performance drops significantly on CULane (only 54.56%). We remark that this highlights the difficulty of launching successful physical backdoor attacks in more diverse and cluttered environments; ii) DBALD achieves the highest ASR across all attack strategies. Our proposed method almost outperforms all baseline methods under LDA, LRA, and LOA. Among these, LDA consistently yields the highest ASR, likely due to its simplicity in altering lane direction. In contrast, LOA and LRA are harder to execute but pose greater safety risks by shifting the vehicle trajectory. LOA achieves higher ASR than LRA, which may be attributed to a broader label shift. Overall, DBALD improves the average ASR by a substantial margin and generalizes well across different lane detection models. This further demonstrates that the physical trigger is more effective in activating the backdoor. iii) DBALD maintains high accuracy on clean samples. Despite its strong attack capability, DBALD introduces only a minor degradation in model performance on clean inputs. Poisoned models retain comparable accuracy to their vanilla counterparts, suggesting that the injected backdoor remains stealthy and does not interfere with normal behavior. iv) ADNet is more robust than other lane detection models. LaneATT, RESA, and SCNN models exhibit high ASRs under DBALD, nearly matching their clean ACCs, indicating high vulnerability. In contrast, ADNet shows greater resilience, with a notable gap of over 16% between its ASR and ACC, suggesting ADNet plays a role in robustness against physical backdoor attacks. More results, e.g., ASR of other triggers and computation resources, are provided in the supplementary.
The Stealthiness of DBALD
To evaluate the stealthiness of injected backdoors, we adopt three representative image-based forensic detection methods due to the lack of LD-specific backdoor detectors. UniDetection (U) [ojha2023towards] uses CLIP-based features to identify subtle inconsistencies in generated content. LGrad (L)[tan2023learning] computes saliency maps from CNN feature gradients to highlight local artifacts. DIRE (D) [wang2023dire] is tailored for detecting diffusion-generated images by modeling visual trace patterns.
We feed both benign and poisoned samples into each detector. All three methods achieve over 90% accuracy on benign samples. The results for poisoned samples are summarized in Table 2, where N/A denotes direct patching methods that require no forensic analysis and are easily detectable by human inspection. BadNets, Blended and LD-Attack remain largely undetected, likely due to their trigger being directly overlaid onto the image without introducing detectable artifacts. In contrast, UniDetection detects nearly 60% of Badlane’s backdoored samples, while only identifying less than 3% of DBALD’s samples. Unfortunately, LGrad fails to effectively detect both DBALD and Badlane, possibly because the detector is not fine-tuned on autonomous driving related scenarios. Additionally, DIRE proves ineffective against DBALD samples. This is expected, as DBALD only modifies a small region of the image while preserving most of the original content. Note that, although BadNets, Blended and LD-Attack exhibit lower detection rates, they are visually conspicuous and thus vulnerable to manual inspection. while the mud trigger from DBALD is visually natural and blends into the road context, making it much harder to identify through manual observation. More trigger visualizations are available in the [dbald_site].
Method CULane TuSimple
U L D U L D
BadNets 2.56 N/A N/A 0.08 N/A N/A
Blended 5.18 N/A N/A 0.17 N/A N/A
LD-Attack 6.27 N/A N/A 1.31 N/A N/A
Badlane 59.31 9.17 N/A 35.60 4.23 N/A
DBALD 2.72 1.76 0.56 0.23 0.51 0.04
Ablation Study
Effectiveness of Heatmap Guidance. We investigate the contribution of heatmap guidance in enhancing the attack robustness of DBALD. All experiments are conducted under the mud trigger on the CULane dataset. Results are given in Table 3, where “Without heatmap” means that the mud trigger is placed randomly during poisoning without any heatmap-based guidance. We find that using heatmap guidance significantly improves ASR on both models: +3.21% on LaneATT and +6.77% on RESA. These results demonstrate that spatially guided trigger placement enhances the robustness of the backdoor attack.
Effectiveness of Diffusion Losses. We further perform ablation on the two proposed diffusion losses, $L_{\text{lane}}$ and $L_{\text{env}}$. Specifically, under forensic detection [ojha2023towards], the detection scores increase from 2.46 (with both losses) to 12.13 (w/o lane loss) and 14.51 (w/o environment loss). These results highlight that both loss terms are essential in producing perceptually natural poisoned images that can better evade existing detection techniques.
Method LaneATT RESA
LDA LRA LOA LDA LRA LOA
With heatmap 81.65 73.45 74.19 78.15 65.15 76.86
Without heatmap 80.97 70.16 68.52 74.23 58.26 69.37
Table 3: Attack performance w/wo the heatmap guidance.
Physical Experiment
To better simulate real-world conditions, we go beyond prior sandbox setups [zhang2024towards] and conduct physical tests in four closed-road environments. We use two types of cones with different shapes and colors to simulate diverse trigger appearances and evaluate attack performance on the poisoned LaneATT model under the LOA setting.
We collect 100 poisoned images using cones across both daytime and nighttime conditions with diverse environments and placements, yielding a 57% ASR. The findings suggest that DBALD remains potentially effective under varying conditions in physical scenarios, and thus poses a real threat to LD models in the physical world. Note that, although lower than simulation-based results, it is expected as the model is not retrained on the physical domain. Additionally, using SOTA weather transfer methods [batifol2025flux, dong2023benchmarking], we apply rain, blur, and occlusion to 100 poisoned images, observing only slight ASR drops (from 57% to 45/48/49%), indicating our method’s robustness to diverse environments. Figure 5 shows benign and attack cases across different cones and lighting conditions.
Figure 5: Physical attacks under different conditions using cone triggers.
Countermeasures
We evaluate several defense strategies against DBALD, following prior works [han2022physical, zhang2024towards]. For fine-tuning, we retrain the poisoned LaneATT model (under LOA) on clean data for 50 epochs, which reduces ASR from 74.19% to 63.68%, showing partial but insufficient mitigation. For neuron pruning, we progressively remove neurons from the final convolutional layer (512 in total) in steps of 50. While pruning 50 neurons only slightly reduces ASR (to 73.67%), more aggressive pruning (e.g., 400 neurons) lowers ASR to 15.87%, but also severely degrades clean accuracy (ACC drops from 74.20% to 21.19%). This suggests pruning is largely ineffective unless it sacrifices benign performance. We also assess a vision-language model Qwen2.5-VL[bai2025qwen3] as an anomaly detector. Sampling 100 poisoned images, we use the model to identify visual anomalies. Detection rates are low: only 6% for cone-based triggers and 3% for mud-based triggers, indicating that multimodal models struggle to detect our triggers. We will investigate the how the quantization and reasoning injection make effect in mitigating the attack for future work [liu2024quantized].
Conclusion and Future Work
We propose DBALD, a diffusion-based backdoor framework for LD systems that places triggers in sensitive regions guided by gradient attention. It supports multiple attack strategies via strategy-specific heatmaps and improves stealthiness through lane preservation and environment consistency losses. Experiments show DBALD outperforms prior backdoors in both digital and physical settings. For future improvements, we believe a task-specific distilled diffusion model could enhance efficiency. While we currently use UltraEdit as a general-purpose backbone, a tailored model trained to seamlessly embed triggers in driving scenes may substantially reduce diffusion steps and costs.
Appendix A Diffusion Denoising
Diffusion-based inpainting methods extend standard diffusion models by introducing spatial constraints, where a mask specifies the region to be edited while preserving the remaining content. Concretely, the model starts from a noisy latent representation and iteratively denoises it, but unlike unconditional generation, the denoising process is conditioned not only on textual instructions but also on the input image and a binary or soft mask that restricts modifications to a target region. In UltraEdit [zhao2024ultraedit], this is implemented via a modified inpainting diffusion pipeline that alternates between global denoising and masked updates, ensuring that edits are localized while maintaining overall image consistency. Building upon this framework, the incorporation of additional losses can be naturally integrated into the denoising process. Therefore, to maintain reality, after decoding the latent into image space two structure-preserving losses constrain the edited region to remain coherent with the original scene. These losses back-propagate through the denoising steps, effectively guiding the noise prediction toward solutions that are both instruction-faithful and visually plausible.
Appendix B More Experiment results
ASR under varying environments
@@ Appendix B More Experiment results
We also find that clean accuracy drops in challenging conditions, which may affect backdoor sensitivity. For instance, under the LOA setting with RESA, clean accuracy decreases from 84.77% (normal) to 82.05% (shadow) and 77.37% (night). This supports our hypothesis that low-light conditions degrade LD model performance, which could inadvertently make the model less reactive to certain perturbations like mud triggers, while still allowing salient triggers (e.g., cones) to remain effective.
@@
ASR under different attack parameters
@@
We further explore the impact of trigger size under the LOA setting using the LaneATT model and CULane dataset with mud triggers. As shown in Figure 6, smaller triggers yield significantly lower attack success rates (ASR), with ASR dropping from 74.19% ($\times40$) to 32.15% ($\times10$). This suggests that tiny triggers are less effective at activating the backdoor. However, ASR slightly drops when the trigger size exceeds $40{\times}40$ (e.g., 70.62% at $50{\times}50$). This is due to a corresponding drop in clean accuracy—excessively large triggers distort the input scene, degrading overall model performance and thus suppressing backdoor activation. Notably, even with a trigger as small as $\times20$ pixels, the ASR exceeds 50%, indicating that DBALD remains effective under compact trigger settings. This supports its potential stealthiness in practical deployment.
ASR under different poisoned ratios
We further study how different poisoning rates affect the attack performance of DBALD. On both LaneATT and RESA models, we train LD models using poisoning rates of 1%, 3%, 5%, 10% and 15% on the CULane dataset. For reference, constructing a 1% poisoned dataset (i.e., 177 poisoned samples) takes approximately 3 minutes. The results are given in Table 5. We find that DBALD achieves relative strong attack performance even at very low poisoning rates. As the poisoning rate rises, the ASR improves gradually, while the accuracy on clean samples decreases slowly.